OpenStudy (anonymous):

((Sin x)/(1-cos x)) + (sin x/(1+cos x)) = 2 csc x

OpenStudy (mertsj):

Are you supposed to prove that is an identity or something?

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

hartnn (hartnn):

ok, so we just need to manipulate the left side to prove it = right side,

OpenStudy (ankit042):

let us start with LHS take the LCM and solve to beign with

hartnn (hartnn):

try this : \(\dfrac{\sin x}{1-\cos x}\dfrac{1+\cos x}{1+\cos x}+\dfrac{\sin x}{1-\cos x}\dfrac{1+\cos x }{1+\cos x}\)

hartnn (hartnn):

now you have a common denominator :) so that you can combine the numerator.

hartnn (hartnn):

what did u get in numerator ?

hartnn (hartnn):

\(\dfrac{\sin x}{1-\cos x}\dfrac{1+\cos x}{1+\cos x}+\dfrac{\sin x}{1+\cos x}\dfrac{1-\cos x }{1-\cos x}\)

hartnn (hartnn):

i corrected it^ did you get it ?

OpenStudy (anonymous):

I haven't done it lol

hartnn (hartnn):

\(\dfrac{\sin x}{1-\cos x}\dfrac{1+\cos x}{1+\cos x}+\dfrac{\sin x}{1+\cos x}\dfrac{1-\cos x }{1-\cos x} \\ = \dfrac{\sin x(1+\cos x)+\sin x(1-\cos x)}{(1-\cos x)(1+\cos x)} \)

OpenStudy (anonymous):

so youre left with sin x ?

hartnn (hartnn):

2 sin x in numerator denominator = \(1-\cos^2 x = \sin^2x\)

OpenStudy (anonymous):

So 1- cos^2 x = sin^2 x

OpenStudy (anonymous):

jk so it's 2 sin x over 1 - cos^2 x = sin^2 x ?

hartnn (hartnn):

yeah, \(\dfrac{2\cancel {\sin x}}{\sin^\cancel{2}x} = \dfrac{2}{\sin x} = 2 \csc x = Right \: \: side\)

OpenStudy (anonymous):

can you write out the full answer, I'm thoroughly confused

hartnn (hartnn):

\(Left \: \: side =\\\dfrac{\sin x}{1-\cos x}\dfrac{1+\cos x}{1+\cos x}+\dfrac{\sin x}{1+\cos x}\dfrac{1-\cos x }{1-\cos x} \\ = \dfrac{\sin x(1+\cos x)+\sin x(1-\cos x)}{(1-\cos x)(1+\cos x)} \\ = \dfrac{\sin x +\sin x \cos x +\sin x - \sin x \cos x}{1-\cos^2x} \\ =\dfrac{2\sin x }{\sin^2x} =\dfrac{2\cancel {\sin x}}{\sin^\cancel{2}x} = \dfrac{2}{\sin x} = 2 \csc x = Right \: \: side\)

OpenStudy (anonymous):

Thank you (:

hartnn (hartnn):

welcome ^_^